OpenStudy (anonymous):

Let x = ln(root4(13)) Then x = A ln(B) where B is an integer what is a and b?

OpenStudy (anonymous):

use ln(a^b) = b ln(a)

OpenStudy (anonymous):

ok do i need to take out of the 4root?

OpenStudy (anonymous):

\[\sqrt{x} = x^{1/2}\]

OpenStudy (anonymous):

yes you need to have an expression that look like Aln(B)

OpenStudy (anonymous):

but since its to the 4 root would it be 13^1/4?

OpenStudy (anonymous):

yep that's it

OpenStudy (anonymous):

so a =1/4 b= 13

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to the 1/3?

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

thanx again

OpenStudy (anonymous):

you're welcome. what do you mean by to the 1/3 ?

OpenStudy (anonymous):

i meant 1/4ln13^1/3

OpenStudy (anonymous):

i subtracted 1 from 1/4

OpenStudy (anonymous):

1/4-1

OpenStudy (anonymous):

1/4*ln 13^1/3 = 1/4 *1/3 * ln 13 = (1/12) *ln 13

OpenStudy (anonymous):

(not very sure i m answering you're question ^^)